minitab Can anyone help me with minitab question?   The life spans of car batteries are normally...

minitab

Can anyone help me with minitab question?

  The life spans of car batteries are normally distributed, with a mean of 44 months and a standard deviation of 5 months.

A car battery is selected at random. Find the probability that the life span of the battery is less than 36 months.

What is the shortest life expectancy a car battery can have and still be in the top 5% of life expectancies?
Answer #1

Ans:

Given that

mean=44

standard deviation=5

1)

z=(36-44)/5=-1.6

P(z<-1.6)=normsdist(-1.6)=0.0548

2)

P(Z>=z)=0.05

P(Z<=z)=1-0.05=0.95

z=normsinv(0.95)=1.645

shorted life expentancy to be in top 5%=44+1.645*5=52.225

For Minitab use:

  1. Choose Calc > Probability Distributions > Normal.
  2. Select Inverse cumulative probability.
  3. In Mean, enter 44.
  4. In Standard deviation, enter 5.
  5. In Input column, enter Probability=0.95